OpenStudy (lukecrayonz):

What is the mass of 1 mole of O2? What is the mass of 1 molecule of O2?

OpenStudy (anonymous):

Oxygen has a n atomic weight of 16 amu, ergo a mole of O_2 has a mass of 32g, and a molecular weight of 32 amu.

OpenStudy (lukecrayonz):

So just 32?

OpenStudy (lukecrayonz):

xishem i have a new one for you:D

OpenStudy (lukecrayonz):

What is the density of O2 at STP?

OpenStudy (lukecrayonz):

And of course STP = volume used, 1 mol any gas = 22.4

OpenStudy (anonymous):

d=MM(P/RT) Ideal gas law.

OpenStudy (lukecrayonz):

o.O what.

OpenStudy (anonymous):

1 mole. of O2 = 32g. 1 mole. at STP occupies 22.4L. Density = Mass / Volume. Density of O2 at STP = 32g / 22.4L = 1.43g/L = 1.43kg/m³